OpenStudy (anonymous):

Show that if P(A) = 0.6, P(B)=0.25 and P( A U B) = 0.7 then A and B are independent.

OpenStudy (nikvist):

\[P(AB)=P(A)+P(B)-P(A\cup B)=0.15=P(A)\cdot P(B)\] \[\Rightarrow\quad independent\]

OpenStudy (anonymous):

ok thanks:)